OpenStudy (anonymous):

In a nuclear reaction, the energy released is equal to 5.4 x 1015 joules. Calculate the mass lost in this reaction. (1 J = 1 kg m2/s2)

OpenStudy (apoorvk):

Hmm. Okay, so have you ever heard of Einstein's E=MC^2 equation? Not academically, just like any vague idea, heard or read about it somewhere?

OpenStudy (anonymous):

Yes in math, it was breifly mentioned once. (:

OpenStudy (apoorvk):

Okay. This is not really maths though. Doesn't matter though. So what the scientists in the beginning of the 20th century found out is that when a nuclear fusion or fission occurs, for eg. when a nucleus breaks up, the product matter (that is the daughter nuclei and neutrons of any), the mass of the nucleus before the fission, and the total mass of the products are NOT found to be EXACTLY equal. So, Einstein put forward a paper 1905 that the lost mass is converted into energy 'E', where: \[E=\Delta M C^2\] where deltaM = change in mass, and 'c' = speed of light.

OpenStudy (apoorvk):

This energy is the energy released during a reaction.

OpenStudy (apoorvk):

*a nuclear reaction, rather.

OpenStudy (anonymous):

6.0 x 10-2 kg 4.86 x 10-1 kg 1.8 x 10-1 kg 1.62 x 102 kg -well out of these answers I think that if I understand it correctly that 1.62x102 kg would be the answer. Am I right?

OpenStudy (anonymous):

sorry thats four different answers kind of looks like i mixed them into one.

OpenStudy (apoorvk):

Let me check. One moment.

OpenStudy (anonymous):

okay

OpenStudy (apoorvk):

Umm, no.. Your answer isn't correct. See, we have E= 5.4 x 10^(15) given, and speed of light 'c'= 3 x 10^8 m/sec. So, just plug these in "E=(deltaM)C^2, and calculate 'deltaM', which is the mass lost. P.S. - delta just means 'change' so 'deltaM' means the change in mass nothing else.

OpenStudy (anonymous):

I have no idea then. /:

OpenStudy (anonymous):

I dont get this at all.

OpenStudy (apoorvk):

Lol, don't fret, this is kid stuff. I have an equation. E=M * C^2 E= 5.4 x 10^15 c= 3 x 10^8 so, \[\large 5.4 \times 10^{15} = M \times (3 \times 10^8)^2\] Now can you find out 'M'?

OpenStudy (apoorvk):

@Jessica18 - fainted or still alive? :P

OpenStudy (anonymous):

I have the equation I just keep getting an answer that doesnt make sense..

OpenStudy (apoorvk):

How are calculating this? Did you take 'c' *squared*?? \[\large c^2 = (3 \times 10^8)^2 = 9 \times 10^{16}\] Did you do this^?

OpenStudy (apoorvk):

Then divide the whole equation by c^2, you should get, \[\large M = \frac{5.4 \times 10^{15}}{9 \times 10^{16}}\] Can you now find out 'M'?

OpenStudy (apoorvk):

What answer are you getting btw? The same as the earlier one? Try solving this^ one now and reply @jessica18